please help me solve this question. y=-3x & 2x-3y=22. they are to the same problem. I can't find the answer. I have to use substitution. I tried for 2 hours and couldn't find the answer. please help me
Answer by ewatrrr(24785)

 
Hi,
y=-3x &
2x-3y=22
2x - 3(-3x) = 22 |Substituting for y
2x + 9x = 22
11x = 22
x = 2 and y = -6 (-2x)
P(2,-6) is the solution for this system of Eqs (Lines intersect at that point)
CHECKING our answer***
Using 2x-3y=22
4 + 18 = 22